# A conjecture on descents, inversions and the weak order on Coxeter groups

### v2

@@ 4.1. Type $B$ permutations, inversions and standardization
$d_{R}(u_{\mathcal{L}})+d_{R}(u_{\mathcal{R}})\leq d_{R}(u).$
Suppose that there is a descent $ca$ in $u_{\mathcal{L}}$ in nonnegative position. This means that $u$ is of the form $\dots c^{\prime}\dots a^{\prime}\dots$ where $a^{\prime}<c^{\prime}$ are letters of $\mathcal{L}(w)$ which are standardized into $a$ and $c$ in $u_{\mathcal{L}}$. As they are consecutive in $u_{\mathcal{L}}$, this means that all letters between $c^{\prime}$ and $a^{\prime}$ are not in $\mathcal{L}(w)$. Then Proposition 4.24 forces the word between $c^{\prime}$ and $a^{\prime}$ to be empty. Therefore $c^{\prime}a^{\prime}$ is a descent of $u$. Moreover, since $\mathcal{L}(w)$ is a symmetric subset of $[\![\bar{n},n]\!]^{*}$, the $B$ standardization cannot change a negative position into a positive one and the descent $c^{\prime}a^{\prime}$ in $u$ is then in nonnegative position; therefore $c^{\prime}a^{\prime}$ is counted in $d_{R}(u)$.
Now suppose that there is a descent $ca$ in $u_{\mathcal{R}}$ (so in positive position). With similar arguments, using Proposition 4.24, we get that there is a descent $c^{\prime}a^{\prime}$ in $u$ with $a^{\prime},c^{\prime}\in\mathcal{R}(w)$. If $c^{\prime}a^{\prime}$ is in nonnegative position, $c^{\prime}a^{\prime}$ is then counted in $d_{R}(u)$. If $c^{\prime}a^{\prime}$ is in negative position, then $\overline{a^{\prime}}~{}\overline{c^{\prime}}$ is a descent in positive position with letters in $\mathcal{F}(w)$ which is counted in $d_{R}(u)$ and was not counted before. This implies the desires inequality and concludes the proof. ∎
@@ 4.4. A remark on the recursive decomposition
Let $w$ in $W$ such that $n$ is in a negative position in $w$. We define the left letters of $w$, $\mathcal{L}(w)$, the forgotten letters of $w$, $\mathcal{F}(w)$, and the right letters of $w$, $\mathcal{R}(w)$, as
$\mathcal{L}(w)$ :=$:=\{a<n\mid(a,n)\not\in\operatorname{invs}(w)\};$
$\mathcal{F}(w)$ $:=\overline{\mathcal{L}(w)}.$
$\mathcal{R}(w)$ :=$:=\{a\mid a\not\in\mathcal{L}(w)\cup\mathcal{F}(w)\};$
@@
$w_{\mathcal{L}}$ $=\operatorname{std}(w_{|\mathcal{L}(w)}),$ $w_{\mathcal{R}}$ $=\operatorname{std}^{B}(w_{|\mathcal{R}(w)})$
$u_{\mathcal{L}}$ $=\operatorname{std}(u_{|\mathcal{L}(w)}),$ $u_{\mathcal{R}}$ $=\operatorname{std}^{B}(u_{|\mathcal{R}(w)})$
$v_{\mathcal{L}}$ $=\operatorname{std}(v_{|\mathcal{L}(w)}),$ $v_{\mathcal{R}}$ $=\operatorname{std}^{B}(v_{|\mathcal{R}(w)}).$
@@
Consider $w=1\bar{3}62\bar{4}\bar{5}|54\bar{2}\bar{6}3\bar{1}$, then $u=1\bar{3}62\bar{5}\bar{4}|45\bar{2}\bar{6}3\bar{1}$ and $v=\bar{6}\bar{4}\bar{5}\bar{3}\bar{2}\bar{1}|123546$ form a bipartition of $w$. We obtain $\mathcal{L}(w)=\{\bar{3},1\}$, $\mathcal{F}(w)=\{\bar{1},3\}$, and $\mathcal{R}(w)=\{\bar{6},\bar{5},\bar{4},\bar{2},2,4,5,6\}$. This gives
@@
$w_{\mathcal{L}}$ $=\operatorname{std}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1\bar{3}})=21,$ $w_{\mathcal{R}}$ $=\operatorname{std}^{B}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}62\bar{4}\bar{5}|54\bar{2}\bar{6}})=41\bar{2}\bar{3}|32\bar{1}\bar{4},$
$u_{\mathcal{L}}$ $=\operatorname{std}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1\bar{3}})=21,$ $u_{\mathcal{R}}$ $=\operatorname{std}^{B}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}62\bar{5}\bar{4}|45\bar{2}\bar{6}})=41\bar{3}\bar{2}|23\bar{1}\bar{4},$
$v_{\mathcal{L}}$ $=\operatorname{std}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{3}1})=12,$ $v_{\mathcal{R}}$ $=\operatorname{std}^{B}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bar{6}\bar{4}\bar{5}\bar{2}|2546})=\bar{4}\bar{2}\bar{3}\bar{1}|1324.$
@@
Consider $w=65\bar{3}4\bar{1}2|\bar{2}1\bar{4}3\bar{5}\bar{6}$, then $u=\bar{5}6\bar{3}4\bar{2}\bar{1}|12\bar{4}3\bar{6}5$, and $v=5\bar{6}\bar{4}\bar{3}\bar{1}2|\bar{2}1346\bar{5}$ form a bipartition of $w$. Since $n$ is in first position ($\bar{n}$ in last), there are no left letters, nor forgotten letters and $w_{\mathcal{L}}=w$, $u_{\mathcal{L}}=u$, $v_{\mathcal{L}}=v$.

### v3

@@ 4.1. Type $B$-permutations, inversions and standardization
$d_{R}(u_{\overleftarrow{n}})+d_{R}(u_{\overrightarrow{n}})\leq d_{R}(u).$
Suppose that there is a descent $ca$ in $u_{\overleftarrow{n}}$ in nonnegative position. This means that $u$ is of the form $\dots c^{\prime}\dots a^{\prime}\dots$ where $a^{\prime}<c^{\prime}$ are letters of ${\overleftarrow{n}}(w)$ which are standardized into $a$ and $c$ in $u_{\overleftarrow{n}}$. As they are consecutive in $u_{\overleftarrow{n}}$, this means that all letters between $c^{\prime}$ and $a^{\prime}$ are not in ${\overleftarrow{n}}(w)$. Then Proposition 4.24 forces the word between $c^{\prime}$ and $a^{\prime}$ to be empty. Therefore $c^{\prime}a^{\prime}$ is a descent of $u$. Moreover, since ${\overleftarrow{n}}(w)$ is a symmetric subset of $\bar{n},n]\!]^{*}$, the $B$ standardization cannot change a negative position into a positive one and the descent $c^{\prime}a^{\prime}$ in $u$ is then in nonnegative position; therefore $c^{\prime}a^{\prime}$ is counted in $d_{R}(u)$.
Now suppose that there is a descent $ca$ in $u_{\overrightarrow{n}}$ (so in positive position). With similar arguments, using Proposition 4.24, we get that there is a descent $c^{\prime}a^{\prime}$ in $u$ with $a^{\prime},c^{\prime}\in{\overrightarrow{n}}(w)$. If $c^{\prime}a^{\prime}$ is in nonnegative position, $c^{\prime}a^{\prime}$ is then counted in $d_{R}(u)$. If $c^{\prime}a^{\prime}$ is in negative position, then $\overline{a^{\prime}}~\overline{c^{\prime}}$ is a descent in positive position with letters in $\mathcal{F}(w)$ which is counted in $d_{R}(u)$ and was not counted before. This implies the desires inequality and concludes the proof. ∎
@@ 4.4. A remark on the recursive decomposition
Let $w$ in $W$ such that $n$ is in a negative position in $w$. We define the left letters of $w$, ${\overleftarrow{n}}(w)$, the forgotten letters of $w$, $\mathcal{F}(w)$, and the right letters of $w$, ${\overrightarrow{n}}(w)$, as
${\overleftarrow{n}}(w)$ $:=\{a<n\mid(a,n)\not\in\operatorname{Inv}(w)\};$
$\mathcal{F}(w)$ $:=\overline{{\overleftarrow{n}}(w)}.$
${\overrightarrow{n}}(w)$ $:=\{a\mid a\not\in{\overleftarrow{n}}(w)\cup\mathcal{F}(w)\};$
@@
$w_{\overleftarrow{n}}$ $=\operatorname{std}(w_{|{\overleftarrow{n}}(w)}),$ $w_{\overrightarrow{n}}$ $=\operatorname{std}^{B}(w_{|{\overrightarrow{n}}(w)})$
$u_{\overleftarrow{n}}$ $=\operatorname{std}(u_{|{\overleftarrow{n}}(w)}),$ $u_{\overrightarrow{n}}$ $=\operatorname{std}^{B}(u_{|{\overrightarrow{n}}(w)})$
$v_{\overleftarrow{n}}$ $=\operatorname{std}(v_{|{\overleftarrow{n}}(w)}),$ $v_{\overrightarrow{n}}$ $=\operatorname{std}^{B}(v_{|{\overrightarrow{n}}(w)}).$
@@
Consider $w=1\bar{3}62\bar{4}\bar{5}|54\bar{2}\bar{6}3\bar{1}$, then $u=1\bar{3}62\bar{5}\bar{4}|45\bar{2}\bar{6}3\bar{1}$ and $v=\bar{6}\bar{4}\bar{5}\bar{3}\bar{2}\bar{1}|123546$ form a bipartition of $w$. We obtain ${\overleftarrow{n}}(w)=\{\bar{3},1\}$, $\mathcal{F}(w)=\{\bar{1},3\}$, and ${\overrightarrow{n}}(w)=\{\bar{6},\bar{5},\bar{4},\bar{2},2,4,5,6\}$. This gives
@@
$w_{\overleftarrow{n}}$ $=\operatorname{std}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1\bar{3}})=21,$ $w_{\overrightarrow{n}}$ $=\operatorname{std}^{B}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}62\bar{4}\bar{5}|54\bar{2}\bar{6}})=41\bar{2}\bar{3}|32\bar{1}\bar{4},$
$u_{\overleftarrow{n}}$ $=\operatorname{std}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1\bar{3}})=21,$ $u_{\overrightarrow{n}}$ $=\operatorname{std}^{B}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}62\bar{5}\bar{4}|45\bar{2}\bar{6}})=41\bar{3}\bar{2}|23\bar{1}\bar{4},$
$v_{\overleftarrow{n}}$ $=\operatorname{std}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{3}1})=12,$ $v_{\overrightarrow{n}}$ $=\operatorname{std}^{B}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bar{6}\bar{4}\bar{5}\bar{2}|2546})=\bar{4}\bar{2}\bar{3}\bar{1}|1324.$
@@
Consider $w=65\bar{3}4\bar{1}2|\bar{2}1\bar{4}3\bar{5}\bar{6}$, then $u=\bar{5}6\bar{3}4\bar{2}\bar{1}|12\bar{4}3\bar{6}5$, and $v=5\bar{6}\bar{4}\bar{3}\bar{1}2|\bar{2}1346\bar{5}$ form a bipartition of $w$. Since $n$ is in first position ($\bar{n}$ in last), there are no left letters, nor forgotten letters and $w_{\overleftarrow{n}}=w$, $u_{\overleftarrow{n}}=u$, $v_{\overleftarrow{n}}=v$.